OpenStudy (anonymous):

A solid aluminum cube measuring 15 inches on a side is melted and rolled into an aluminum wire that is 0.2 inches in diameter. What will be the length of this wire?

OpenStudy (amistre64):

prolly .6cms and then the machine gets clooged and breaks down

OpenStudy (amistre64):

15^3 = 3375

OpenStudy (amistre64):

3375/pi (.01) = length

OpenStudy (amistre64):

107429.5866 ??

OpenStudy (anonymous):

im trying hang on... LOL My machine been busted!

OpenStudy (amistre64):

:) the wire is a tiny column; base area * length = area of cube

OpenStudy (amistre64):

.2/2 = .1 ; .1(.1) = .01 .01 pi

OpenStudy (amistre64):

.01pi * L = 15^3 L = 3375/.01pi = 337500/pi

OpenStudy (amistre64):

337 500 / pi = 107,429.587...

OpenStudy (anonymous):

approximately 107,429.6 inches That is no tiny piece of wire... Our answers match!! Thank you

OpenStudy (anonymous):

Hey up top you got 3375/pi(.01) you get 107... I got 10.742958...

OpenStudy (anonymous):

can you see if I made a mistake